OpenStudy (anonymous):

Find the volume that remains after a hole of radius 1 is bored through the center of a solid cylinder of radius 2 and height 4

OpenStudy (tkhunny):

Cylinder Volume is \(\pi\cdot 2^{2}\cdot 4\) Hole Volume is \(\pi\cdot 1^{2}\cdot 4\)